Anuradha Sharma, a start up entrepreneur from Bareilly, has invested Rs 80 lacs in an apparel retail

store. Business has been good, and the store shows an accounting profit of Rs 10 lacs for the last year.

This profit is after taxes and after payment of a Rs 20 lacs salary to Ms. Sharma. This salary is less than

what she could make at another job, which is about equal to Rs 40 lacs. Considering the risk involved in

the fashion retail business post Covid’19, she believes that a 15 percent after-tax rate of return is

appropriate for this type of investment. (20 marks)

a. Given this information, calculate the economic profit earned by Ms Sharma.

b. What accounting profit would the firm have to earn in order for the firm to break even in term

of economic profit?


1
Expert's answer
2021-12-03T08:36:26-0500

a. The economic profit earned by Ms Sharma is:

"EP = 10 - (40 + 0.15\u00d780) = -42."

b. The firm would have to earn the accounting profit of 10 + 42 = 52 in order for the firm to break even in term of economic profit.